### v5

@@ 4.8 Effectiveness of Large Language Model (RQ4)
The baselines’ performance on the LRT task is extremely poor, highlighting the need for improvement (RQ2). When we attempted a more refined approach using the ST models ($\mathcal{K}$ashif), it performed better than the baselines but did not achieve satisfactory results on an unseen dataset (RQ3). This indicates that the ST models can partially address some of the issues inherent in the baseline approaches. However, the ST models lack the robustness needed to generalize effectively across unseen data, as discussed in RQ3. Given their promising results across many tasks AroraHH24 ; abs_2310_18648 , RQ4 aims to assess whether prompting LLMs with an engineered prompt offers a meaningful alternative for LRT and to compare its effectiveness against prompts that were not engineered. We posit that LLMs, with their pre-training across different domains, might significantly improve trace link recovery.
Metholodology. As discussed in Section 3.3, we designed a prompt, based on the Rice structure vogelsang2024using . We prompted the GPT-4o model to generate recommendations for trace links between requirements and GDPR provisions. We base our analysis on the four documents discussed in RQ3. We compare the LLM’s recommendations, generated with our prompt for each requirement, against our ground truth. We begin by evaluating Rice_LRT under two temperature values (0 and 0.5) and five different random seed values. Then, we compare its performance to the baseline prompts using their reported best settings.
@@
In contrast, increasing randomness by setting the temperature equal to 0.5 caused a drop in F2 score for a constant seed value. The highest drop was observed for seed 16, where the F2 score dropped from 61.1% to 56.9%. Moreover, we observe that the model’s behavior is not consistent across test documents. For example, when using seed value 13374224, increasing the temperature from 0 to 0.5 results in a decrease in F2 score for RD1, RD2, and RD4, while it leads to an improvement for RD3. Nonetheless, we observe that Rice_LRT shows little variation across different seeds when the temperature is set to 0, suggesting that random variation in results is primarily due to higher temperature. Thus, we select the performance using temperature 0 and arbitrarily select seed 16 to compare against the baselines.
@@
Table 9 shows the results of the Rice_LRT approach and the baseline prompts. Rice_LRT outperforms all prompts at both the trace and requirement levels. P1 includes a retrieval phase that selects the top-k artifacts based on semantic similarity. In our case, the artifacts are the 26 GDPR provisions. Therefore, we experimented with two k values: 13 and 26, corresponding to half and all the GDPR provisions, respectively. Since using k=13 resulted in a 12.6% loss of provisions during the retrieval step by the semantic similarity technique, we chose to include all provisions, given that the total number of provisions is not excessively large in our case. P1 exhibits a lower number of retrieved links (TP+FP) compared to Rice_LRT and P2. Based on our observations of the rationale generated by the model, this stems primarily from its inability to capture indirect relationships between requirements and their corresponding provisions. For instance, a requirement stating that a user’s profile must include at least one charging point may seem purely functional but implicitly involves handling personal data, such as user identity or location. Consequently, it should be linked to the GDPR provision on obtaining user consent for data processing. The prompt’s inability to capture this implicit connection results in lower retrieval performance and the omission of true trace links. Moreover, P1’s F2 score is only 0.8 pp higher than $\mathcal{K}$ashif (13.5% in Table 7), which suggests that this prompt does not fully leverage GPT-4o’s capabilities, while $\mathcal{K}$ashif already captures much of the underlying knowledge by being tuned over another dataset, allowing it to remain highly competitive. P2 exhibits better performance, achieving a higher retrieval compared to P1. We observed that the rationales generated by P2 provide more details in explaining how to link specific requirements to the relevant provisions. While P2’s high retrieval rate increased the number of TPs, it concurrently increased the FPs. This ultimately led to a higher F2 score than P1. P3_1 is similar to P2 in terms of a low retrieval rate, while P3_2 resulted in producing the highest FPs compared to all other prompt templates. Moreover, $\mathcal{K}$ashif has a higher recall compared to P3_1 and P3_2 has a lower F2 score compared to $\mathcal{K}$ashif. On the other hand, the inclusion of context and detailed instructions in Rice_LRT yields more effective outputs from GPT-4o, leading to improved performance at both the trace link and requirement levels. It is worth noting that Rice_LRT, under all temperature and random seed configurations reported in Table 8, consistently outperforms the baselines. Nonetheless, an investigation into which specific elements of the prompt design influence the model’s behavior falls outside the scope of the current study.
@@
At the requirements level, the number of exact matches is very low for Rice_LRT. We note that Rice_LRT outputs at least one regulatory code for each requirement (based on our prompt of Section 3.3) even when requirements do not have any trace links in the ground truth. This is one explanation for the sharp decrease in exact matches. Despite this, the number of partial matches has increased significantly, thereby improving the overall success rate. While one would ideally like an approach with a high exact-match rate, we note that the results remain beneficial, as we discuss next.
@@
Fig. 7 shows the split of partially matched requirements for the number of FPs. For instance, for RD1, there were 56 partially matched requirements. Of these, 24 (42.9%) had only one FP, 25 (44.6%) had two FPs, and the remaining 7 (12.5%) had three FPs. As seen in the figure, across all four documents, there were very few requirements with four FPs. This indicates that most partially matched requirements had a manageable number of FPs, typically ranging from one to three. This result is significant because it suggests that the model’s outputs are not overwhelming for analysts to process. Fewer FPs per requirement allow analysts to efficiently review and validate the suggested trace links, reducing their cognitive load. Instead of starting from scratch or sifting through a vast set of 26 possible provisions per requirement, analysts can focus on validating and refining a much smaller, pre-filtered set of trace links. This aligns with the principle of assisted decision-making skitka1999does , in which automated tools augment human judgment by narrowing options.
@@
Our results further indicate that the GPT-4o model successfully demonstrated a stronger understanding of the LRT task when provided with an engineered prompt. This indicates that Rice_LRT is effective at identifying the underlying logic and rationale behind provisions, even with a limited number of few-shot examples. Its ability to navigate complex relationships and extract logical links demonstrates its robustness in understanding the nuances of regulatory requirements. However, the cases it misses highlight areas where the connections may require deeper domain-specific knowledge or additional context to resolve ambiguities.
@@
On investigating the FPs for each requirement, we observed that several predicted trace links may be relevant depending on the application context, even though they do not exactly match the ground truth. These false positives include provisions that are not in the ground truth but are relevant to the input requirements, or provisions in the ground truth (but not all of them), which are why they are considered partial matches. This underscores the potential of Rice_LRT to identify trace links that correspond to potential associations between requirements and provisions that may not have been contemplated when building the ground truth. Such cases could still be informative to the analysts. For example, the Rice_LRT output presented in Section 3.3 included three predictions with corresponding rationales. Of these, [SEC] is the ground truth, and [ACC] and [CNF] are categorized as FPs. The rationale for [ACC] highlights that requiring a key file ensures proper authentication, which can be interpreted as supporting the right to access. Similarly, the rationale for [CNF] emphasizes that protecting the database with a key file ensures sensitive data remains confidential. While these codes are not explicitly part of the ground truth for this requirement, they surface related regulatory considerations that may enrich the analyst’s understanding of the requirement and its broader implications under GDPR. Hence, while FPs may not align perfectly with the ground truth, their contextual relevance, as inferred from the generated rationale, can offer valuable insights for the LRT task. This also underscores the inherent subjectivity of the LRT task, especially when dealing with broadly framed regulations like GDPR, which often leave room for interpretation, compared to domain-specific regulations such as HIPAA.
@@
To assess the statistical significance of the differences in proportions of TPs versus FNs (recall) and TPs versus FPs (precision) between $\mathcal{K}$ashif and Rice_LRT, we employ Fisher’s exact test fisher1922interpretation . This test calculates the probability of obtaining the observed data (TPs, FPs, and FNs) under the null hypothesis that proportions are equal across techniques. The p-values for RD1 for the proportions of (TPs, FPs) and (TPs, FNs) are 3e-4 and 1e-5, respectively. For RD2, the corresponding values are 2.5e-2 and 1e-5, for RD3, 4.6e-3 and 1e-5, and for RD4, 0 and 1e-5. The comparison between Rice_LRT and P1, , P3_1, and P3_2 yields values comparable to those reported for $\mathcal{K}$ashif, while the results for P2 are presented as follows: 0 and 5e-1, 1e-4 and 6e-1, 1e-4 and 5e-2, and 1e-4 and 3e-2 for RD1, RD2, RD3, and RD4, respectively. As we can see, across the four test documents, the resulting p-value is very small (below 5e-2). It falls well below the usual significance levels, indicating that the differences in precision and recall among $\mathcal{K}$ashif, P1, P2, P3_1, , P3_2 and Rice_LRT are statistically significant in most cases, except for P2 and the recall scores for RD1 to RD2.
The answer to RQ4 is that our Rice-based approach, which utilizes an engineered prompting on GPT-4o, outperforms the baselines on the LRT task across the four test documents. Further, the engineered context and instructions used in the few-shot setting of Rice_LRT achieved the highest F2 score by maximizing TPs while reducing FPs compared to the baselines. It also significantly outperforms $\mathcal{K}$ashif, although $\mathcal{K}$ashif remains competitive against P1 and P3_2 prompts.
@@ 4.9 Discussion
We experimented with a variety of classifiers, including two IR-, one ML-, two LM-, and one LLM-based methods. We compared them against our ST-based classifier, $\mathcal{K}$ashif, on the LRT task using the publicly available HIPAA dataset. The results show that $\mathcal{K}$ashif achieves the highest F2 score among all baselines, including B, TraceBERT, and LLaMA, which have been previously proposed for inner-requirements traceability. Moreover, $\mathcal{K}$ashif achieved a strong overall ranking performance, obtaining an MAP of 81% and the highest discrimination across thresholds, with an AUC of 0.93. $\mathcal{K}$ashif demonstrated a superior ability to address the linguistic discrepancies between requirements and provisions, outperforming existing models in the literature that target traceability across different artifact types. We also demonstrated that Rice_LRT, as an engineered prompt, surpasses baseline prompts introduced in prior work, achieving a recall of 84%, a precision of $\approx$30%, and F2 score of 61% on a newly created, more complex dataset using the GDPR provisions. From a more practical standpoint, we analyzed Rice_LRT ’s performance per requirement, using both Macro Recall (MR) and Cost. Rice_LRT achieves an MR of 84% with a Cost of 9.5%, meaning that, on average, only 9.5% of the total provisions are retrieved, while ensuring that 84% of the true trace links are included in the retrieved list. This represents a 36 pp improvement in MR over the best baseline prompt, while reducing Cost by 13 pp. Moreover, we show that simple Yes/No–style prompts achieve low MR and Cost retrieval on the test documents, indicating that they do not fully leverage GPT-4o’s capabilities. Lastly, our $\mathcal{K}$ashif classifier achieved competitive F2 scores on our newly created GDPR dataset compared to P1 and P3_2 prompts, despite being fine-tuned only on the HIPAA dataset.
@@ 5 Threats to Validity
Internal Validity. Bias is a well-known internal validity concern. To mitigate bias, in RQ3 and RQ4, the dataset of over four documents was curated by two annotators with more than a decade of experience in RE. Before the traceability sessions, there was no exposure to the technical details of our approach. The second potential threat to internal validity concerns the few-shot prompting in RQ4. The initial few-shot examples used for GPT-4o’s prompt engineering could introduce confirmation bias, potentially influencing the model’s predictions. To mitigate this, we designed the few-shot examples to reflect realistic usage scenarios where LLM is a recommendation tool guided by a human expert’s rationale for the first few requirements. This approach aligns with practical applications while minimizing the risk of confirmation bias. Additionally, the limited number of examples in the few-shot prompt was deliberately chosen to avoid overfitting. By doing so, we allowed the LLM sufficient flexibility to independently apply reasoning across the remaining requirements, maintaining a balance between guidance and adaptability. This approach ensures the LLM’s outputs remain broadly applicable while minimizing potential validity threats, as seen by the relatively high success rate in RQ4.
External Validity. We evaluated $\mathcal{K}$ashif on two datasets, namely HIPAA and four new documents against GDPR. HIPAA is a pre-existing dataset frequently used in the RE literature. The test documents used in RQ3 and RQ4 (four new documents against GDPR), which we created as part of our work, covers two types of textual requirements, including user stories and shall-type requirements. Such diversity helped increase the generalizability of our results. Experiments on more diverse requirements documents and other regulations are nonetheless required to improve the external validity of our study.

### v6

@@ 4.8 Effectiveness of Large Language Model (RQ4)
The baselines’ performance on the LRT task is extremely poor, highlighting the need for improvement (RQ2). When we adopted a more refined approach using the ST models ($\mathcal{K}$ashif), it outperformed the baselines but did not yield satisfactory results on an unseen dataset (RQ3). This indicates that ST models can partially address issues inherent in baseline approaches. However, the ST models lack the robustness needed to generalize effectively across unseen data, as discussed in RQ3. Given their promising results across many tasks AroraHH24 ; abs_2310_18648 , RQ4 aims to assess whether prompting LLMs with an engineered prompt offers a meaningful alternative for LRT and to compare its effectiveness against prompts that were not engineered. We posit that LLMs, with their pretraining across diverse domains, may significantly improve trace link recovery.
Metholodology. As discussed in Section 3.3, we designed a prompt, based on the Rice structure vogelsang2024using . We prompted the GPT-4o model to generate recommendations for trace links between requirements and GDPR provisions. We base our analysis on the four documents discussed in RQ3. We compare the LLM’s recommendations, generated with our prompt for each requirement, against our ground truth. We begin by evaluating Rice_LRT under two temperature values (0 and 0.5) and five different random seed values. We then compare its performance with the baseline prompts using their reported best settings.
@@
In contrast, increasing randomness by setting the temperature equal to 0.5 caused a drop in F2 score for a constant seed value. The largest drop was observed for seed 16, from 61.1% to 56.9%. Moreover, we observe that the model’s behavior is not consistent across test documents. For example, when using a seed value of 13374224, increasing the temperature from 0 to 0.5 decreases the F2 score for RD1, RD2, and RD4, whereas it increases it for RD3. Nonetheless, we observe that Rice_LRT shows little variation across different seeds when the temperature is set to 0, suggesting that random variation in results is primarily due to higher temperature. Thus, we select the performance at temperature 0 and arbitrarily set the seed to 16 to compare against the baselines.
Moreover, we evaluated Rice_LRT using different number of few-shot examples (3, 3, and 4) to assess the sensitivity of its performance to the degree of few-shot supervision. We randomly chose 2, 3, and 4 examples from the existing few-shot examples. This created three different few-shot sets for each document. We ran the model separately with each set and recorded the resulting F2 scores. Based on the results, the average F2 score across 2-, 3-, and 4-shot settings are 56.5% (-4.6 pp), 58.9% (-2.2 pp), and 59.5% (-1.6 pp), respectively. Nonetheless, Rice_LRT consistently outperforms the baseline approaches across all example counts.
@@
Table 9 shows the results of the Rice_LRT approach and the baseline prompts. Rice_LRT outperforms all prompts at both the trace and requirement levels. P1 includes a retrieval phase that selects the top-k artifacts based on semantic similarity. In our case, the artifacts are the 26 GDPR provisions. Therefore, we experimented with two k values: 13 and 26, corresponding to half and all the GDPR provisions, respectively. Since using k=13 resulted in a 12.6% loss of provisions during the retrieval step by the semantic similarity technique, we chose to include all provisions, given that the total number of provisions is not excessively large in our case. P1 exhibits a lower number of retrieved links (TP+FP) compared to Rice_LRT and P2. Based on our observations of the model’s rationale, this stems primarily from its inability to capture indirect relationships between requirements and their corresponding provisions. For instance, a requirement stating that a user’s profile must include at least one charging point may seem purely functional but implicitly involves handling personal data, such as user identity or location. Consequently, it should be linked to the GDPR provision governing the obtaining of user consent for data processing. The prompt’s inability to capture this implicit connection results in lower retrieval performance and the omission of true trace links. Moreover, P1’s F2 score is only 0.8 pp higher than $\mathcal{K}$ashif (13.5% in Table 7), which suggests that this prompt does not fully leverage GPT-4o’s capabilities, while $\mathcal{K}$ashif already captures much of the underlying knowledge by being tuned over another dataset, allowing it to remain highly competitive. P2 exhibits better performance, achieving a higher retrieval compared to P1. We observed that the rationales generated by P2 provide more detail on how to link specific requirements to the relevant provisions. While P2’s high retrieval rate increased the number of TPs, it concurrently increased the FPs. This ultimately led to a higher F2 score than P1. P3_1 is similar to P2 in terms of a low retrieval rate, while P3_2 resulted in producing the highest FPs compared to all other prompt templates. Moreover, $\mathcal{K}$ashif has a higher recall compared to P3_1, and P3_2 has a lower F2 score compared to $\mathcal{K}$ashif. On the other hand, the inclusion of context and detailed instructions in Rice_LRT yields more effective outputs from GPT-4o, leading to improved performance at both the trace link and requirement levels. It is worth noting that Rice_LRT, under all temperature and random seed configurations reported in Table 8, consistently outperforms the baselines. Nonetheless, an investigation into which specific elements of the prompt design influence the model’s behavior falls outside the scope of the current study.
@@
At the requirements level, the number of exact matches is very low for Rice_LRT. We note that Rice_LRT outputs at least one regulatory code for each requirement (based on our prompt of Section 3.3) even when requirements do not have any trace links in the ground truth. This is one explanation for the sharp decrease in exact matches. Despite this, the number of partial matches has increased significantly, thereby improving the overall success rate. While an approach with a high exact-match rate would be ideal, the results remain beneficial, as discussed next.
@@
Fig. 7 shows the split of partially matched requirements for the number of FPs. For instance, for RD1, there were 56 partially matched requirements. Of these, 24 (42.9%) had only one FP, 25 (44.6%) had two FPs, and the remaining 7 (12.5%) had three FPs. As seen in the figure, across all four documents, there were very few requirements with four FPs. This indicates that most partially matched requirements had a manageable number of FPs, typically between 1 and 3. This result is significant because it suggests that the model’s outputs are not overwhelming for analysts to process. Fewer FPs per requirement allow analysts to review and validate the proposed trace links more efficiently, reducing their cognitive load. Instead of starting from scratch or sifting through a vast set of 26 possible provisions per requirement, analysts can focus on validating and refining a much smaller, pre-filtered set of trace links. This aligns with the principle of assisted decision-making skitka1999does , in which automated tools augment human judgment by narrowing the set of options.
@@
Our results further indicate that the GPT-4o model successfully demonstrated a stronger understanding of the LRT task when provided with an engineered prompt. This indicates that Rice_LRT is effective at identifying the underlying logic and rationale behind provisions, even with a limited number of few-shot examples. Its ability to navigate complex relationships and extract logical links demonstrates its robustness in understanding the nuances of regulatory requirements. However, the cases it misses highlight areas in which the connections may require deeper, domain-specific knowledge or additional context to resolve ambiguities.
@@
On investigating the FPs for each requirement, we observed that several predicted trace links may be relevant depending on the application context, even though they do not exactly match the ground truth. These false positives include provisions that are not in the ground truth but are relevant to the input requirements, as well as provisions in the ground truth (though not all of them), which is why they are considered partial matches. This underscores the potential of Rice_LRT to identify trace links that indicate associations between requirements and provisions that may not have been considered during ground-truth construction. Such cases could still be informative to the analysts. For example, the Rice_LRT output presented in Section 3.3 included three predictions with corresponding rationales. Of these, [SEC] is the ground truth, and [ACC] and [CNF] are categorized as FPs. The rationale for [ACC] highlights that requiring a key file ensures proper authentication, which can be interpreted as supporting the right to access. Similarly, the rationale for [CNF] emphasizes that protecting the database with a key file ensures sensitive data remains confidential. While these codes are not explicitly part of the ground truth for this requirement, they surface related regulatory considerations that may enrich the analyst’s understanding of the requirement and its broader implications under GDPR. Hence, while FPs may not align perfectly with the ground truth, their contextual relevance, as inferred from the generated rationale, can offer valuable insights for the LRT task. This also underscores the inherent subjectivity of the LRT task, especially when dealing with broadly framed regulations like GDPR, which often leave room for interpretation, compared to domain-specific regulations such as HIPAA.
@@
To assess the statistical significance of the differences in proportions of TPs versus FNs (recall) and TPs versus FPs (precision) between $\mathcal{K}$ashif and Rice_LRT, we employ Fisher’s exact test fisher1922interpretation . This test calculates the probability of obtaining the observed data (TPs, FPs, and FNs) under the null hypothesis that proportions are equal across techniques. The p-values for RD1 for the proportions of (TPs, FPs) and (TPs, FNs) are 3e-4 and 1e-5, respectively. For RD2, the corresponding values are 2.5e-2 and 1e-5, for RD3, 4.6e-3 and 1e-5, and for RD4, 0 and 1e-5. The comparison between Rice_LRT and P1, , P3_1, and P3_2 yields values comparable to those reported for $\mathcal{K}$ashif, while the results for P2 are presented as follows: 0 and 5e-1, 1e-4 and 6e-1, 1e-4 and 5e-2, and 1e-4 and 3e-2 for RD1, RD2, RD3, and RD4, respectively. Across the four test documents, the resulting p-value is very small (below 5e-2). It falls well below the usual significance levels, indicating that the differences in precision and recall among $\mathcal{K}$ashif, P1, P2, P3_1, P3_2, and Rice_LRT are statistically significant in most cases, except for P2 and the recall scores for RD1 to RD2.
The answer to RQ4 is that our Rice-based approach, which utilizes an engineered prompting on GPT-4o, outperforms the baselines on the LRT task across the four test documents. Furthermore, the engineered context and instructions used in the few-shot setting of Rice_LRT achieved the highest F2 score by maximizing TPs while minimizing FPs relative to the baselines. It also significantly outperforms $\mathcal{K}$ashif, although $\mathcal{K}$ashif remains competitive against P1 and P3_2 prompts.
@@ 4.9 Discussion
We experimented with a variety of classifiers, including two IR-, one ML-, two LM-, and one LLM-based methods. We compared them against our ST-based classifier, $\mathcal{K}$ashif, on the LRT task using the publicly available HIPAA dataset. The results show that $\mathcal{K}$ashif achieves the highest F2 score among all baselines, including B, TraceBERT, and LLaMA, which have been previously proposed for inter-requirements traceability. Moreover, $\mathcal{K}$ashif achieved a strong overall ranking performance, obtaining an MAP of 81% and the highest discrimination across thresholds, with an AUC of 0.93. $\mathcal{K}$ashif demonstrated a superior ability to address the linguistic discrepancies between requirements and provisions, outperforming existing models in the literature that target traceability across different artifact types. We also demonstrated that Rice_LRT, as an engineered prompt, surpasses baseline prompts introduced in prior work, achieving a recall of 84%, a precision of $\approx$30%, and F2 score of 61% on a newly created, more complex dataset using the GDPR provisions. From a more practical standpoint, we analyzed Rice_LRT ’s performance per requirement, using both Macro Recall (MR) and Cost. Rice_LRT achieves an MR of 84% with a Cost of 9.5%, meaning that, on average, only 9.5% of the total provisions are retrieved, while ensuring that 84% of the true trace links are included in the retrieved list. This represents a 36-pp improvement in MR over the best baseline prompt, while reducing Cost by 13 pp. Moreover, we show that simple Yes/No–style prompts achieve low MR and Cost retrieval on the test documents, indicating that they do not fully leverage GPT-4o’s capabilities. Lastly, our $\mathcal{K}$ashif classifier achieved competitive F2 scores on our newly created GDPR dataset compared to P1 and P3_2 prompts, despite being fine-tuned only on the HIPAA dataset.
@@ 5 Threats to Validity
Internal Validity. Bias is a well-known internal validity concern. To mitigate bias, for RQ3 and RQ4, the dataset comprising more than four documents was curated by two annotators with more than a decade of experience in RE. Before the traceability sessions, there was no exposure to the technical details of our approach. The second potential threat to internal validity concerns the few-shot prompting in RQ4. The initial few-shot examples used for GPT-4o’s prompt engineering could introduce confirmation bias, potentially influencing the model’s predictions. To mitigate this, we designed the few-shot examples to reflect realistic usage scenarios in which an LLM serves as a recommendation tool guided by a human expert’s rationale for the first few requirements. This approach aligns with practical applications while minimizing the risk of confirmation bias. Additionally, the limited number of examples in the few-shot prompt was deliberately chosen to avoid overfitting. By doing so, we provided the LLM sufficient flexibility to independently apply reasoning to the remaining requirements, while maintaining a balance between guidance and adaptability. This approach ensures the LLM’s outputs remain broadly applicable while minimizing potential validity threats, as seen by the relatively high success rate in RQ4.
External Validity. We evaluated $\mathcal{K}$ashif on two datasets, namely HIPAA and four new documents against GDPR. HIPAA is a pre-existing dataset frequently used in the RE literature. The test documents used in RQ3 and RQ4 (four new documents against GDPR), which we created as part of our work, cover two types of textual requirements, including user stories and shall-type requirements. Such diversity helped increase the generalizability of our results. Experiments on more diverse requirements documents and other regulations are nonetheless required to improve the external validity of our study. Moreover, our evaluation is based on GPT-4o since this model is widely used in requirements engineering. Results obtained with other LLMs may differ; however, we expect them to be similar to those of LLMs with the same capabilities (focused on textual data and reasoning), and we leave a multi-model evaluation to future work.